# SSR-Encoder: Encoding Selective Subject Representation for Subject-Driven Generation

### 26 Dec 2023 v1

@@ Abstract
1Shanghai Jiao Tong University, 2Xiaohongshu Inc.,
3Beijing University Of Posts And Telecommunications, 4Carnegie Mellon University, 5ShanghaiTech University
@@
11footnotetext: Work done during internship at Xiaohongshu Inc.22footnotetext: Equally contributed.33footnotetext: Corresponding author.
@@ 2 Related Work
Text-to-image diffusion models. In recent years, text-to-image diffusion models [36, 37, 39, 41, 42, 34, 50, 2, 32, 53] has made remarkable progress, particularly with the advent of diffusion models, which have propelled text-to-image generation to large-scale commercialization. DALLE [36] first achieved stunning image generation results using an autoregressive model. Subsequently, DALLE2 [37] employed a diffusion model as the generative model, further enhancing text-to-image synthesis ability. Imagen [41] and Stable Diffusion [39] trained diffusion models on larger datasets, further advancing the development of diffusion models and becoming the mainstream for image generation large models. DeepFloyd IF [42] utilized a triple-cascade diffusion model, significantly enhancing the text-to-image generation capability, and even generating correct fonts. Stable Diffusion XL [34], a two-stage cascade diffusion model, is the latest optimized version of stable diffusion, greatly improving the generation of high-frequency details, small object features, and overall image color.
@@ 3.2 Subject Conditioned Generation
$\begin{split}\mathit{O}&=\underbrace{\operatorname{CrossAttntion}\left(\mathbf{Q},\mathbf{K},\mathbf{V},\mathit{c_{t}},\mathit{x_{t}}\right)}_{\textrm{text condition}}\\ &+\lambda\underbrace{\operatorname{CrossAttntion}\left(\mathbf{Q},\mathbf{K_{S}},\mathbf{V_{S}},\mathit{c_{s}^{\prime}},\mathit{x_{t}}\right)}_{\textrm{subject condition}},\end{split}$ (4)
where $\mathit{c_{t}}$ represents the text embedding and $\mathit{x_{t}}$ represents the latent. $\mathbf{Q},\mathbf{K},\mathbf{V}$ represents query, key, and value projection layers in the original text branch respectively while $\mathbf{K_{S}},\mathbf{V_{S}}$ represents trainable copies of key and value projection layers for concatenated subject embedding $\mathit{c_{s}}$. We set $\lambda$ to 1 in all experimental settings if not otherwise specified.
@@ 4.1 Experimental Setup
Implementation details. We employed Stable Diffusion V1-5 as the pre-trained diffusion model, complemented by the pre-trained CLIP text encoder. For training, images are resized to ensure the shortest side is 512 pixels, followed by a center crop to achieve a 512$\times$512 resolution, and sent to the stable diffusion. The same image is resized to 224$\times$224 and sent to the SSR encoder. The model underwent 1,000,000 iterations of training on 8 H800s GPUs, with a batch size of 16 per GPU and a learning rate of 1e-4. Inference was performed using DDIM as the sampler, with a step size of 30 and a guidance scale set to 7.5.
@@ 4.2 Evaluation Metrics
For a comprehensive comparison with state-of-the-art (SOTA) methods, we employed the following metrics: DINO Scores[33] and DINO-M Scores to assess subject alignment, CLIP-T [16] for evaluating image-text alignment, CLIP Exclusive Score (CLIP-ES) to measure the exclusivity of subject representation, and the Aesthetic Score [43] to gauge the overall quality of the generated images.
@@ 4.4 Experiment Results
Table 2: Quantitative comparison of different methods. Metrics that are bold and underlined represent methods that rank 1st and 2nd, respectively. Our approach outperforms other finetuning-free methods, and offers competitive results among the finetuning-based ones.
Type Method CLIP-T $\uparrow$ CLIP-ES $\downarrow$ DINO-M $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ Aesthetic Score $\uparrow$
(Multi-subject bench) (DreamBench)
Textual Inversion 0.240 0.212 0.410 0.255 0.569 6.029
Finetune-based Dreambooth 0.298 0.223 0.681 0.305 0.668 6.330
methods Break-A-Scene 0.285 0.187 0.630 0.287 0.653 6.234
@@
BLIP-Diffusion 0.287 0.198 0.514 0.300 0.594 6.212
Reference only 0.242 0.195 0.434 0.187 0.503 5.812
Finetune-free IP-adapter 0.272 0.201 0.442 0.283 0.574 6.432
methods ELITE 0.253 0.194 0.483 0.232 0.561 6.283
@@ 5 Conclusion
Supplementary
Appendix A Preliminaries
A.1 Preliminary for Diffusion Models
@@ A.2 Preliminary for CLIP
Appendix B Designing Choice of Image Encoder
In this section, we conduct a preliminary reconstruction experiment to demonstrate that vanilla image features fail to capture fine-grained representations of the target subject and verify the effectiveness of our method. We first introduce our experimental setup and evaluation metrics in Sec. B.1. Subsequently, we explain the implementation details of each setting in Sec. B.2. Finally, we conduct qualitative and quantitative experiments in Sec. B.3 to prove the superiority of our proposed methods compared to previous works.
B.1 Experimental Setup
@@ B.3 Experiment Results
Appendix C Details of Comparison Experiments
C.1 Details of Compared Methods
@@ C.2 Details of Implementation
Appendix D User Study
We conducted a user study to compare our method with DB, TI, Break-A-Scene, ELITE, and IP-adapter perceptually. For each evaluation, each user will see one input image with multiple concepts, two different prompts for different concepts, and 5 images generated by each prompt and each method. 60 evaluators were asked to rank each generated image from 1 (worst) to 5 (best) concerning its selectivity, text-image alignment, subject alignment, and generative quality. The results are shown in Table. LABEL:fig:secB_userstudy indicate that our method outperforms the comparison methods in generative quality and better balances subject consistency and text-image alignment.
Appendix E Human Image Generation
@@
Appendix F Examples of Evaluation Samples
@@
Appendix G Details of Our Training Data and the Multi-subject Bench
@@
Appendix H Compatibility with ControlNet
@@
Appendix I Compatibility with AnimateDiff

### 14 Mar 2024 v2

@@ Abstract
1Shanghai Jiao Tong University, 2Xiaohongshu Inc., 3Beijing University of Posts and Telecommunications,
4National University of Singapore, 5Carnegie Mellon University, 6ShanghaiTech University
@@
11footnotetext: Work done during internship at Xiaohongshu Inc.22footnotetext: Corresponding authors.
@@ 2 Related Work
Text-to-image diffusion models. In recent years, text-to-image diffusion models [38, 39, 40, 42, 43, 36, 51, 2, 34, 54] have made remarkable progress, particularly with the advent of diffusion models, which have propelled text-to-image generation to large-scale commercialization. DALLE [38] first achieved stunning image generation results using an autoregressive model. Subsequently, DALLE2 [39] employed a diffusion model as the generative model, further enhancing text-to-image synthesis ability. Imagen [42] and Stable Diffusion [40] trained diffusion models on larger datasets, further advancing the development of diffusion models and becoming the mainstream for image generation large models. DeepFloyd IF [43] utilized a triple-cascade diffusion model, significantly enhancing the text-to-image generation capability, and even generating correct fonts. Stable Diffusion XL [36], a two-stage cascade diffusion model, is the latest optimized version of stable diffusion, greatly improving the generation of high-frequency details, small object features, and overall image color.
@@ 3.2 Subject Conditioned Generation
$\begin{split}\mathit{O}&=\underbrace{\operatorname{CrossAttention}\left(\mathbf{Q},\mathbf{K},\mathbf{V},\mathit{c_{t}},\mathit{x_{t}}\right)}_{\textrm{text condition}}\\ &+\lambda\underbrace{\operatorname{CrossAttention}\left(\mathbf{Q},\mathbf{K_{S}},\mathbf{V_{S}},\mathit{c_{s}^{\prime}},\mathit{x_{t}}\right)}_{\textrm{subject condition}},\end{split}$ (4)
where $\mathit{c_{t}}$ represents the text embedding and $\mathit{x_{t}}$ represents the latent. $\mathbf{Q},\mathbf{K},\mathbf{V}$ represents query, key, and value projection layers in the original text branch respectively while $\mathbf{K_{S}},\mathbf{V_{S}}$ represents trainable copies of key and value projection layers for concatenated subject embedding $\mathit{c_{s}}$. $\lambda$ is a weight adjustment factor, with a default value of 1.
@@ 4.1 Experimental Setup
Implementation details. We employed Stable Diffusion V1-5 as the pre-trained diffusion model, complemented by the pre-trained CLIP text encoder. For training, images are resized to ensure the shortest side is 512 pixels, followed by a center crop to achieve a 512$\times$512 resolution, and sent to the stable diffusion. The same image is resized to 224$\times$224 and sent to the SSR encoder. The model training process is divided into two steps. In the first step, the multi-scale strategy is not employed, and the model is trained for 1 million steps on 8H800s GPUs, with a batch size of 16 per GPU and a learning rate of 5e-5. In the second step, the same hyper-parameters are used, and the model parameters obtained from the first step are used as the initialization parameters. The multi-scale strategy is employed in this step to train the model for an additional 100,000 steps. Inference was performed using DDIM as the sampler, with a step size of 30 and a guidance scale set to 7.5.
@@ 4.2 Evaluation Metrics
For a comprehensive comparison with state-of-the-art (SOTA) methods, we employed the following metrics: DINO Scores[4], CLIP-I[37] and DINO-M Scores to assess subject alignment, CLIP-T [18] for evaluating image-text alignment, CLIP Exclusive Score (CLIP-ES) to measure the exclusivity of subject representation, and the Aesthetic Score [44] to gauge the overall quality of the generated images.
@@ 4.4 Experiment Results
Table 2: Quantitative comparison of different methods. Metrics that are bold and underlined represent methods that rank 1st and 2nd, respectively. ${}^{\dagger}$ indicates that the experimental value is referenced from BLIP-Diffusion[26].
Type Method CLIP-T $\uparrow$ CLIP-ES $\downarrow$ DINO-M $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$ Aesthetic
@@
Textual Inversion 0.240 0.212 0.410 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ 6.029
Finetune-based Dreambooth 0.298 0.223 0.681 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ 6.330
methods Break-A-Scene 0.285 0.187 0.630 0.287 0.653 0.788 6.234
[…]
BLIP-Diffusion 0.287 0.198 0.514 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ 6.212
Reference only 0.242 0.195 0.434 0.286 0.542 0.727 5.812
Finetune-free IP-adapter 0.272 0.201 0.442 0.274 0.608 0.809 6.432
methods ELITE 0.253 0.194 0.483 0.298 0.605 0.775 6.283
@@ 5 Conclusion
Supplementary Material
In this supplementary material, we first introduce the preliminaries of Diffusion and CLIP in Section F. Following that, we provide an in-depth discussion on our Detail-Preserving Image Encoder in Section G. In subsequent sections, we introduce the methods we compared against and the user study we conducted, specifically in Section H and Section I respectively. We also present our results on human image generation in Section J. Additional results from our work on Dreambench and Multi-subject bench are showcased in Section L. We then provide further details about our training data and Multi-subject bench in Section M. In Section N and Section O, we present the outcomes generated by combining our SSR-encoder with ControlNet [55] and animatediff [15], which not only demonstrates the generalization of our SSR encoder but also illustrates its seamless applicability in the realm of controllable generation and video generation for maintaining character consistency with reference images. Lastly, we analyze the broader impact brought by our method and the limitation of our method in Section P and Section Q.
F Preliminaries
F.1 Preliminary for Diffusion Models
@@ F.2 Preliminary for CLIP
G Designing Choice of Image Encoder
In this section, we conduct a preliminary reconstruction experiment to demonstrate that vanilla image features fail to capture fine-grained representations of the target subject and verify the effectiveness of our method. We first introduce our experimental setup and evaluation metrics in Sec. G.1. Subsequently, we explain the implementation details of each setting in Sec. G.2. Finally, we conduct qualitative and quantitative experiments in Sec. G.3 to prove the superiority of our proposed methods compared to previous works.
G.1 Experimental Setup
@@ G.3 Experiment Results
H Details of Comparison Experiments
H.1 Details of Compared Methods
@@ H.2 Details of Implementation
I User Study
We conducted a user study to compare our method with DB, TI, Break-A-Scene, ELITE, and IP-adapter perceptually. For each evaluation, each user will see one input image with multiple concepts, two different prompts for different concepts, and 5 images generated by each prompt and each method. 60 evaluators were asked to rank each generated image from 1 (worst) to 5 (best) concerning its selectivity, text-image alignment, subject alignment, and generative quality. The results are shown in Table. 9 indicate that our method outperforms the comparison methods in generative quality and better balances subject consistency and text-image alignment.
J Human Image Generation
@@ I User Study
K Ablations of $\tau$ and $\lambda$
As shown in Fig. 10 (a), under the same training settings, when $\tau$ was 0.01, the model managed to balance both identity consistency and selectivity. The effects of different $\lambda$ values on the images under ablation and fixed seed conditions are shown in Fig. 10 (b). The smaller $\lambda$, the weaker the influence of the reference image.
Figure 10: Visual ablation results of $\tau$ and $\lambda$.
L Examples of Evaluation Samples
@@
M Details of Our Training Data and the Multi-subject Bench
@@
N Compatibility with ControlNet
@@
O Compatibility with AnimateDiff